# Molecule Graph Networks with Many-body Equivariant Interactions

### v2

@@ Appendix A Mathematical Foundations and Analysis of the ENINet Architecture
Let $\rho:{\rm O}(3)\hookrightarrow{\rm GL}(3)$ be the canonical group representation of ${\rm O}(3)$, let $\varphi=\pi_{d}\oplus\rho:{\rm O}(3)\rightarrow{\rm GL}(\mathbb{R}^{d}\oplus\mathbb{R}^{3})$, and let $\psi:{\rm O}(3)\rightarrow{\rm GL}(\mathbb{R}^{d}\otimes\mathbb{R}^{3})$ defined in Definition S11. In particular, the function $\phi_{j}$ is the composition of the mappings
@@
where the first mapping is $(\varphi,\varphi)$-equivariant (by Example S10 and Proposition S5), and the second mapping is $(\varphi,\psi)$-equivariant (by Example S11). By Proposition S4, the function $\phi_{j}$ is ${\rm O}(3)$-equivariant with respect to the representations $\varphi$ and $\psi$.
For a given graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ and a vertex $i\in\mathcal{V}$, the ENINet architecture aggregates directional information from the neighboring vertices $j\in\mathcal{N}(i)$, where $\mathcal{N}(i)$ denotes the set of neighbors of $i$. This aggregation process is guided by a core function that encodes the relative positional and feature information between $i$ and its neighbors, facilitating the representation of local geometric and structural properties of the graph. Specifically, the core function is defined by
$\theta:\bigoplus_{j\in\mathcal{N}(i)}\mathbb{R}^{d}\oplus\mathbb{R}^{3}\longrightarrow\mathbb{R}^{d}\otimes\mathbb{R}^{3},\ \ ((\vec{\mathbf{f}}_{ji},\vec{\textbf{r}}_{ji}))_{j\in\mathcal{N}(i)}\longmapsto\sum_{j\in\mathcal{N}(i)}w_{ji}\cdot\left(\vec{\mathbf{f}}_{ji}\otimes\frac{\vec{\textbf{r}}_{ji}}{\|\vec{\textbf{r}}_{ji}\|}\right),$ (46)
@@
$\bigoplus_{j\in\mathcal{N}(i)}\mathbb{R}^{d}\oplus\mathbb{R}^{3}\xrightarrow{\ \ \phi\ \ }\bigoplus_{j\in\mathcal{N}(i)}\mathbb{R}^{d}\otimes\mathbb{R}^{3}\xrightarrow{\ \ \omega\ \ }\bigoplus_{j\in\mathcal{N}(i)}\mathbb{R}^{d}\otimes\mathbb{R}^{3}\xrightarrow{\ \ \sigma\ \ }\mathbb{R}^{d}\otimes\mathbb{R}^{3}$ (47)
with group representations $\bigoplus_{j\in\mathcal{N}(i)}\varphi$, $\bigoplus_{j\in\mathcal{N}(i)}\psi$, $\bigoplus_{j\in\mathcal{N}(i)}\psi$, and $\psi$, where the functions $\phi$, $\omega$, and $\sigma$ are defined as in the following equation:
$\begin{split}\phi\left(((\vec{\mathbf{f}}_{ji},\vec{\textbf{r}}_{ji}))_{j\in\mathcal{N}(i)}\right)&=(\phi_{j}(\vec{\mathbf{f}}_{ji},\vec{\textbf{r}}_{ji}))_{j\in\mathcal{N}(i)},\\ \omega\left((\mathfrak{v}_{j})_{j\in\mathcal{N}(i)}\right)&=(w_{ji}\cdot\mathfrak{v}_{j})_{j\in\mathcal{N}(i)},\\ \sigma\left((\mathfrak{v}_{j})_{j\in\mathcal{N}(i)}\right)&=\sum_{j\in\mathcal{N}(i)}\mathfrak{v}_{j}.\end{split}$
In particular, since $\phi_{j}$ is ${\rm O}(3)$-equivariant with respect to the representations $\varphi$ and $\psi$ for each $j\in\mathcal{N}(i)$, it follows from Proposition S5 that the function $\phi$ is also ${\rm O}(3)$-equivariant with respect to the corresponding group representations. On the other hand, the following lemmas (Lemma S1 and Lemma S2) show that $\omega$ and $\sigma$ are ${\rm O}(3)$-equivariant with respect to the corresponding group representations.
@@
since $Q\in{\rm GL}(n)$ is $\mathbb{R}$-linear. In particular, $f$ is ${\rm O}(n)$-equivariant with respect to $\varphi$ and $\rho$. ∎
@@
By combining the results of Lemmas S1 and S2 with the decomposition (47), the core function $\theta$ defined in Equation (46) is shown to be ${\rm O}(3)$-equivariant with respect to the corresponding representations. This result is formally stated as Theorem S1.
@@
The $2$-body core aggregation function defined in Equation (46) is ${\rm O}(3)$-equivariant.
For the $3$-body feature aggregation, the ENINet leverages the line graph of the underlying graph to incorporate directional information related to $3$-body interactions. Specifically, for a graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ and a vertex $i\in\mathcal{V}$ with distinct neighbors $j$ and $k$, the 3-body feature aggregation framework of ENINet additionally incorporates the directional vector
@@
$\sum_{j\in\mathcal{N}(i)}w_{ji}\cdot\left(\vec{\mathbf{f}}_{ji}\otimes\frac{\vec{\textbf{r}}_{ji}}{\|\vec{\textbf{r}}_{ji}\|}\right)+\sum_{j,k\in\mathcal{N}(i),\ j\neq k}w_{jk}\cdot\left(\vec{\mathbf{f}}_{jk}\otimes\frac{\vec{\textbf{r}}_{jk}}{\|\vec{\textbf{r}}_{jk}\|}\right),$ (48)
@@
$\theta^{(3)}:\left(\bigoplus_{j\in\mathcal{N}(i)}\mathbb{R}^{d}\oplus\mathbb{R}^{3}\right)\oplus\left(\bigoplus_{j,k\in\mathcal{N}(i),j\neq k}\mathbb{R}^{d}\oplus\mathbb{R}^{3}\right)\longrightarrow(\mathbb{R}^{d}\otimes\mathbb{R}^{3})\oplus(\mathbb{R}^{d}\otimes\mathbb{R}^{3}),$ (49)
where $\theta^{(3)}$ maps each summand in the domain to the corresponding summand in the range space. Using similar arguments as those employed to prove the equivariance of $\theta$, along with Proposition S5, it follows that the function $\theta^{(3)}$ is also ${\rm O}(3)$-equivariant. This result is formally stated in the following theorem.
@@
The $3$-body core aggregation function defined in Equation (49) is ${\rm O}(3)$-equivariant.
More generally, incorporating the line graphs and embeddings defined in Section A.1 allows for the consideration of $N$-body feature aggregation. Specifically, for a graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, the sequence of line graphs $\mathcal{G}^{(0)}=\mathcal{G},\mathcal{G}^{(1)},\mathcal{G}^{(2)},\dots$ is obtained iteratively. Furthermore, given the coordinate embedding $\nu^{(0)}:\mathcal{V}^{(0)}\rightarrow\mathbb{R}^{3}$, the embedding functions $\nu^{(i)}:\mathcal{V}^{(i)}\rightarrow\mathbb{R}^{3}$ for $i=1,2,\dots,N$ are defined. By iteratively proceeding with the aggregation process, as in the cases of $\theta=\theta^{(2)}$ and $\theta^{(3)}$, the core aggregation function $\theta^{(N)}$ is constructed.
By Proposition S1, transforming the coordinates using a matrix $Q\in{\rm O}(3)$ also transforms the higher-dimensional line graph coordinate embeddings $\nu^{(n)}:\mathcal{V}^{(n)}\rightarrow\mathbb{R}^{3}$ by multiplying the matrix $Q$. That is, the coordinate embeddings of the line graphs are ${\rm O}(3)$-equivariant. By combining the results of Theorems S1, S2, and the previous findings, the ${\rm O}(3)$-equivariance property can be extended to the case of $N$-body feature aggregations:
@@
The $N$-body core aggregation function defined above, with the coordinate embedding functions $\nu^{(i)}:\mathcal{V}^{(i)}\rightarrow\mathbb{R}^{3}$ for $i=0,1,\dots,N$ (cf. Definition S3), is ${\rm O}(3)$-equivariant.
In summary, Section A.5 focuses on analyzing the ENINet architecture’s feature aggregation mechanism using ${\rm O}(3)$-equivariant functions. Starting with $2$-body feature aggregation, it defines directional edge features and their aggregation via the core function $\theta$, demonstrating its ${\rm O}(3)$-equivariance through a decomposition framework. This is extended to $3$-body interactions using line graphs, incorporating additional directional vectors to define a “higher-dimensional” core function $\theta^{(3)}$, also shown to be ${\rm O}(3)$-equivariant. The section concludes by generalizing this approach to $N$-body interactions.

### v3

@@ Appendix A Mathematical Foundations and Analysis of the ENINet Architecture
Let $\rho:\operatorname{O}(3)\hookrightarrow{\rm GL}(3)$ be the canonical group representation of $\operatorname{O}(3)$, let $\varphi=\pi_{d}\oplus\rho:\operatorname{O}(3)\rightarrow{\rm GL}(\mathbb{R}^{d}\oplus\mathbb{R}^{3})$, and let $\psi:\operatorname{O}(3)\rightarrow{\rm GL}(\mathbb{R}^{d}\otimes\mathbb{R}^{3})$ defined in Definition S11. In particular, the function $\phi_{j}$ is the composition of the mappings
@@
where the first mapping is $(\varphi,\varphi)$-equivariant (by Example S10 and Proposition S5), and the second mapping is $(\varphi,\psi)$-equivariant (by Example S11). By Proposition S4, the function $\phi_{j}$ is $\operatorname{O}(3)$-equivariant with respect to the representations $\varphi$ and $\psi$.
For a given graph $\mathscr{G}=(\mathscr{V},\mathscr{E})$ and a vertex $i\in\mathscr{V}$, the ENINet architecture aggregates directional information from the neighboring vertices $j\in\mathscr{N}(i)$, where $\mathscr{N}(i)$ denotes the set of neighbors of $i$. This aggregation process is guided by a core function that encodes the relative positional and feature information between $i$ and its neighbors, facilitating the representation of local geometric and structural properties of the graph. Specifically, the core function is defined by
$\theta:\bigoplus_{j\in\mathscr{N}(i)}\mathbb{R}^{d}\oplus\mathbb{R}^{3}\longrightarrow\mathbb{R}^{d}\otimes\mathbb{R}^{3},\ \ ((\vec{\mathbf{f}}_{ji},\vec{\textbf{r}}_{ji}))_{j\in\mathscr{N}(i)}\longmapsto\sum_{j\in\mathscr{N}(i)}w_{ji}\cdot\left(\vec{\mathbf{f}}_{ji}\otimes\frac{\vec{\textbf{r}}_{ji}}{\|\vec{\textbf{r}}_{ji}\|}\right),$ (48)
@@
$\bigoplus_{j\in\mathscr{N}(i)}\mathbb{R}^{d}\oplus\mathbb{R}^{3}\xrightarrow{\ \ \phi\ \ }\bigoplus_{j\in\mathscr{N}(i)}\mathbb{R}^{d}\otimes\mathbb{R}^{3}\xrightarrow{\ \ \omega\ \ }\bigoplus_{j\in\mathscr{N}(i)}\mathbb{R}^{d}\otimes\mathbb{R}^{3}\xrightarrow{\ \ \sigma\ \ }\mathbb{R}^{d}\otimes\mathbb{R}^{3}$ (49)
with group representations $\bigoplus_{j\in\mathscr{N}(i)}\varphi$, $\bigoplus_{j\in\mathscr{N}(i)}\psi$, $\bigoplus_{j\in\mathscr{N}(i)}\psi$, and $\psi$, where the functions $\phi$, $\omega$, and $\sigma$ are defined as in the following equation:
$\begin{split}\phi\left(((\vec{\mathbf{f}}_{ji},\vec{\textbf{r}}_{ji}))_{j\in\mathscr{N}(i)}\right)&=(\phi_{j}(\vec{\mathbf{f}}_{ji},\vec{\textbf{r}}_{ji}))_{j\in\mathscr{N}(i)},\\ \omega\left((\mathfrak{v}_{j})_{j\in\mathscr{N}(i)}\right)&=(w_{ji}\cdot\mathfrak{v}_{j})_{j\in\mathscr{N}(i)},\\ \sigma\left((\mathfrak{v}_{j})_{j\in\mathscr{N}(i)}\right)&=\sum_{j\in\mathscr{N}(i)}\mathfrak{v}_{j}.\end{split}$
In particular, since $\phi_{j}$ is $\operatorname{O}(3)$-equivariant with respect to the representations $\varphi$ and $\psi$ for each $j\in\mathscr{N}(i)$, it follows from Proposition S5 that the function $\phi$ is also $\operatorname{O}(3)$-equivariant with respect to the corresponding group representations. On the other hand, the following lemmas (Lemma S1 and Lemma S2) show that $\omega$ and $\sigma$ are $\operatorname{O}(3)$-equivariant with respect to the corresponding group representations.
@@
since $\rho_{g}:W\rightarrow W$ is $\mathbb{R}$-linear. In particular, $\sigma$ is $G$-equivariant with respect to $\varphi$ and $\rho$. ∎
@@
By combining the results of Lemmas S1 and S2 with the decomposition (49), the core function $\theta$ defined in Equation (48) is shown to be $\operatorname{O}(3)$-equivariant with respect to the corresponding representations. This result is formally stated as Theorem S1.
@@
The 222-body core aggregation function defined in Equation (48) is $\operatorname{O}(3)$-equivariant with respect to the representations induced by the canonical representation $\operatorname{O}(3)\hookrightarrow\operatorname{GL}(3)$.
For the 333-body feature aggregation, the ENINet leverages the line graph of the underlying graph to incorporate directional information related to 333-body interactions. Specifically, for a graph $\mathscr{G}=(\mathscr{V},\mathscr{E})$ and a vertex $i\in\mathscr{V}$ with distinct neighbors $j$ and $k$, the 3-body feature aggregation framework of ENINet additionally incorporates the directional vector
@@
$\sum_{j\in\mathscr{N}(i)}w_{ji}\cdot\left(\vec{\mathbf{f}}_{ji}\otimes\frac{\vec{\textbf{r}}_{ji}}{\|\vec{\textbf{r}}_{ji}\|}\right)+\sum_{j,k\in\mathscr{N}(i),\ j\neq k}w_{jk}\cdot\left(\vec{\mathbf{f}}_{jk}\otimes\frac{\vec{\textbf{r}}_{jk}}{\|\vec{\textbf{r}}_{jk}\|}\right),$ (50)
@@
$\theta^{(3)}:\left(\bigoplus_{j\in\mathscr{N}(i)}\mathbb{R}^{d}\oplus\mathbb{R}^{3}\right)\oplus\left(\bigoplus_{j,k\in\mathscr{N}(i),j\neq k}\mathbb{R}^{d}\oplus\mathbb{R}^{3}\right)\longrightarrow(\mathbb{R}^{d}\otimes\mathbb{R}^{3})\oplus(\mathbb{R}^{d}\otimes\mathbb{R}^{3}),$ (51)
where $\theta^{(3)}$ maps each summand in the domain to the corresponding summand in the range space. Using similar arguments as those employed to prove the equivariance of $\theta$, along with Proposition S5, it follows that the function $\theta^{(3)}$ is also $\operatorname{O}(3)$-equivariant. This result is formally stated in the following theorem.
@@
The 333-body core aggregation function defined in Equation (51) is $\operatorname{O}(3)$-equivariant with respect to the representations induced by the canonical representation $\operatorname{O}(3)\hookrightarrow\operatorname{GL}(3)$.
More generally, incorporating the line graphs and embeddings defined in Section A.1 allows for the consideration of $N$-body feature aggregation. Specifically, for a graph $\mathscr{G}=(\mathscr{V},\mathscr{E})$, the sequence of line graphs $\mathscr{G}^{(0)}=\mathscr{G},\mathscr{G}^{(1)},\mathscr{G}^{(2)},\dots$ is obtained iteratively. Furthermore, given the coordinate embedding $\nu^{(0)}:\mathscr{V}^{(0)}\rightarrow\mathbb{R}^{3}$, the embedding functions $\nu^{(i)}:\mathscr{V}^{(i)}\rightarrow\mathbb{R}^{3}$ for $i=1,2,\dots,N$ are defined. By iteratively proceeding with the aggregation process, as in the cases of $\theta=\theta^{(2)}$ and $\theta^{(3)}$, the core function $\theta^{(N)}$ for the $N$-body feature aggregation is constructed.
By Proposition S1, transforming the coordinates using a matrix $Q\in\operatorname{O}(3)$ also transforms the higher-dimensional line graph coordinate embeddings $\nu^{(n)}:\mathscr{V}^{(n)}\rightarrow\mathbb{R}^{3}$ by multiplying the matrix $Q$. That is, the coordinate embeddings of the line graphs are $\operatorname{O}(3)$-equivariant. By combining the results of Theorems S1, S2, and the previous findings, the $\operatorname{O}(3)$-equivariance property can be extended to the case of $N$-body feature aggregations:
@@
The $N$-body core aggregation function defined above, with the coordinate embedding functions $\nu^{(i)}:\mathscr{V}^{(i)}\rightarrow\mathbb{R}^{3}$ for $i=0,1,\dots,N$ (cf. Definition S3), is $\operatorname{O}(3)$-equivariant with respect to the representations induced by the canonical representation $\operatorname{O}(3)\hookrightarrow\operatorname{GL}(3)$.
In summary, Section A.5 focuses on analyzing the ENINet architecture’s feature aggregation mechanism using $\operatorname{O}(3)$-equivariant functions. Starting with 222-body feature aggregation, it defines directional edge features and their aggregation via the core function $\theta$, demonstrating its $\operatorname{O}(3)$-equivariance through a decomposition framework. This is extended to 333-body interactions using line graphs, incorporating additional directional vectors to define a “higher-dimensional” core function $\theta^{(3)}$, also shown to be $\operatorname{O}(3)$-equivariant. The section concludes by generalizing this approach to $N$-body interactions.